OpenStudy (anonymous):

3y squared-12y/9y cubed

OpenStudy (agent47):

(3y^2-12y)/9y = 3y(y-4)/9y = (y-4)/3

jhonyy9 (jhonyy9):

so do you think 3y^2 -12y ---------- = ? 9y^3 yes ?

jhonyy9 (jhonyy9):

are you here ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

how is this exersice ? how i have wrote there ?

jhonyy9 (jhonyy9):

do you need simplifie it ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so in numerator what can you factorizing out ?